OpenStudy (anonymous):

What is the new volume of a rectangular prism that is originally 12 cubic meters and is then resized using a scale factor of 3 on each side length? a. 324 ^3 b. 4 m^3 c. 108^3 d. 36m^3 Please help !

OpenStudy (anonymous):

Ratios when talking about volume are cube. a : b = 3 : 1 For volume, a^2 : b^2 = 27 : 1 Therefore, set up a proportion. 27/1 = x/12 x = 27(12) x = 324

OpenStudy (anonymous):

thanks (:

OpenStudy (anonymous):

You're welcome :)